### v3

Abstract
[Abstract] Technical debt (TD) represents the long-term costs associated with suboptimal design or code decisions in software development, often made to meet short-term delivery goals. Self-Admitted Technical Debt (SATD) occurs when developers explicitly acknowledge these trade-offs in the codebase, typically through comments or annotations. Automated detection of SATD has become an increasingly important research area, particularly with the rise of natural language processing (NLP), machine learning (ML), and deep learning (DL) techniques that aim to streamline SATD detection.
This systematic literature review provides a comprehensive analysis of SATD detection approaches published between 2014 and 2024, focusing on the evolution of techniques from NLP-based models to more advanced ML, DL, and Transformers-based models such as BERT. The review identifies key trends in SATD detection methodologies and tools, evaluates the effectiveness of different approaches using metrics like precision, recall, and F1-score, and highlights the primary challenges in this domain, including dataset heterogeneity, model generalizability, and the explainability of models.
The findings suggest that while early NLP methods laid the foundation for SATD detection, more recent advancements in DL and Transformers models have significantly improved detection accuracy. However, challenges remain in scaling these models for broader industrial use. This SLR offers insights into current research gaps and provides directions for future work, aiming to improve the robustness and practicality of SATD detection tools.
[…]
keywords:
self-admitted technical debt, SATD, detection, identification, categorization, systematic review
articletype: Review Articlejournal: Journalvolume: 00
1 Introduction
Business needs often require developers to be able to deliver new products and features as rapidly as possible, in order to reach the market. In fact, a developer must be able to balance carefully designing applications and seamlessly coding their features. In the business or corporate world, however, schedules and limited resources frequently hinder developers from generating perfect code before deploying it into products 1. This becomes commonplace when stakeholders cut several technical processes to catch up on momentum, especially in startup practices 2, 3. A tight deadline frequently leads to technical debt, as developers lack time to consider long-term solutions 4.
Technical debt (TD) is often used to describe the trade-off that occurs when an organization (or a developer) chooses to prioritize the delivery of new features or functionality over the long-term maintenance of the software 5. TD can occur in any software development project and can be caused by a variety of factors, including lack of resources, time, knowledge, or experience on the part of the development team 6, 7.
Among various types of TD, self-admitted technical debt (SATD) has gained considerable attention due to its explicit acknowledgment within the source code comments by developers. SATD represents instances where developers intentionally or unintentionally leave behind suboptimal design or implementation choices, often with a clear understanding of the resultant negative impact on the codebase 8. This acknowledgment is not merely an admission of suboptimal practices but a vital piece of information that can guide future development activities. Recognizing SATD is crucial for maintaining software quality and sustainability, as it can accumulate over time and impede future development efforts.
Over the past decade, many approaches to detecting SATD have emerged, driven by advances in natural language processing (NLP), machine learning (ML), and deep learning (DL). These approaches aim to automate the detection of SATD, facilitate their management, and helping software teams address it proactively. Alongside these approaches, various methodologies have been proposed to detect SATD in different stages of the software development life cycle, leveraging manual and semi-automated, and automated techniques. However, the diversity of techniques, methodologies, tools, and evaluation metrics used in SATD detection has led to a fragmented landscape, making it challenging for researchers and practitioners to gain a holistic understanding of the field.
To address this fragmentation, we selected a study that aligns closely with our focus on SATD detection. We summarize this work, compare it with our systematic literature review (SLR), and provide a condensed analysis of the key differences. Our SLR distinguishes itself from prior reviews in several key areas:
Time frame and scope.
Our review is the most extensive to date, covering relevant studies published between 2014 and 2024. In contrast, the review by Sierra et al. focuses on a narrower time frame, covering studies published between 2014 and 2018 9.
Systematic approach.
We followed the original and rigorous procedures outlined by Kitchenham and Charters 10 for conducting systematic reviews. In contrast, Sierra et al.’s study 9 appears to be based on a survey approach, without explicitly detailing how studies were sourced.
Comprehensiveness.
In addition to relying on search engines, our review included a bidirectional snowballing process to ensure we captured all relevant articles. As a result, we analyzed a broader range of studies.
Profoundness.
Unlike Sierra et al. 9, who did not report the metric evaluation results, our review includes detailed analysis of metrics such as precision, recall, and F1-score, categorizing them based on the approaches and methodologies used.
Analysis.
While Sierra et al. 9 broadly discussed SATD activities without focusing deeply on detection approaches, our review provides a more detailed evaluation of detection techniques, their effectiveness, and the tools proposed for SATD detection.
In this study, we conduct and evaluate a systematic literature review that focuses on the detection of SATD, particularly using NLP-based approaches and their combinations. The objectives and key contributions of our study are as follows:
  • We consolidate and categorize various approaches employed to detect SATD, revealing various techniques and identifying critical trends over the past decade.
  • We examine methodologies proposed for SATD detection, showcasing their impact on automation and accuracy.
  • We investigate tools developed for SATD detection over the past decade, providing an overview of their effectiveness.
  • We compare the best-performing approaches by evaluating precision, recall, F1-score, and generalizability across different software systems and datasets.
  • We identify primary challenges in SATD detection, offering insights for future research and improvement.
In addition to the Introduction, the remaining study sections are organized as follows: Section 2 introduces the concept and background, and Section  3 describes the methodology utilized to conduct the SLR. Sections 4 and 5 summarize and discuss the obtained results, while Section 6 identifies the threats to validity of this study. Finally, Section 7 offers a conclusion.
2 Background
This section introduces the concepts underpinning this research, related to NLP and self-admitted technical debt, as well as the relevant works.
2.1 Self-Admitted Technical Debt (SATD)
The term “technical debt” was coined by software developer Ward Cunningham. He initially used the metaphor for communicating to non-technical stakeholders the importance of budgeting resources for refactoring 5. Technical debt, sometimes referred to as tech debt or code debt, is the result of a development team rushing the delivery of projects or features that must later be refactored. In other words, there is a trade-off between delivering high-quality, effective, and optimized code and the short-term benefits of achieving release dates 6, 4. A small amount of debt is acceptable because it can increase productivity, but if left uncontrolled, it could end up costing a lot of time, extra work, and money in the future 11, 12.
The term “not-quite-right” is commonly used to refer to internal software development activities that are postponed but may lead to future difficulties if not addressed 6, 13. It represents the debt incurred when development teams opt for quick and easy solutions in the short term, risking long-term negative impacts.
The term “TD” encompasses below-par aspects of the software that developers are aware of but lack resources to address, such as outdated documentation, unperformed tests, complex code needing refactoring, and known defects left unattended 14, 15. These immature artifacts can lead to challenges in meeting quality standards and cause unforeseen delays in implementing necessary changes. Numerous empirical studies have recently concentrated on examining source code comments and other textual artifacts that indicate technical debt, often known as self-admitted technical debt (SATD) 8. The source used to detect SATD is usually in the form of natural language often expressed through text, such as code comments, commit messages, or other textual artifacts; therefore, SATD detection significantly relies on NLP-based approaches.
2.2 Natural Language Processing (NLP)
Natural Language Processing (NLP) is an interdisciplinary field that studies the interaction between computers and human language, encompassing tasks like speech recognition, natural language understanding, and natural language generation 16, 17. Chowdhury 17 explains that the objective of NLP researchers is to understand how people perceive and use language so that the appropriate tools and techniques can be developed to assist computers in understanding and using natural languages to complete the necessary tasks.
NLP deals with processing natural language, which is the language used by humans in text or audio form. Computational linguistics, rule-based human language modeling, statistical models, machine learning, and deep learning are combined to enable computer programs to interpret and process human language 18, 19. In the context of SATD detection, plain text is used as input, including project artifacts such as source code comments, commit messages, pull requests, and issue tracking systems 20, 21.
Source code comments serve as annotations within the code, providing documentation directly incorporated by programmers 22. These comments are essential for understanding a system and may contain important information, such as the purpose of code additions, progress status on collective tasks, or reasons for specific changes 23. Comments can express concerns that need attention in the future (e.g., the TODO keyword), identify emerging issues, and guide future actions 24, 25. They provide valuable context specific to their authors and are human-readable.
Recently, automated approaches utilizing NLP have gained traction in addressing software engineering problems, including the detection of TD, which is called SATD detection. NLP has been effectively used to automatically identify the presence of TD and even categorize specific types of SATD. Notably, numerous studies have integrated NLP with machine learning (ML) models or deep learning (DL) architectures and also Transformers, aiming to enhance model performance (evidence is provided in Subsection 4.1).
3 Research Methodology
This study is carried out in accordance with the methodology of conducting an SLR described by Kitchenham et al26, 27, 10. The procedures for SLR are broken down into three stages: planning, conducting, and reporting the SLR 10. We developed a review protocol during the planning stage that included the following steps: i) identification of research topics, ii) creation of a search strategy, iii) study selection criteria, iv) study quality assessment, v) data extraction procedure, and vi) data synthesis process (as illustrated in Figure  1).
Figure 1: Graphical representation of the SLR process used in this article
The review process comprises several steps. Initially, research questions were formulated for the SLR. Subsequently, a search strategy was devised, involving the identification of search strings and the selection of online databases for locating primary studies. The third stage encompassed the identification of pertinent studies aligned with the research questions, along with the establishment of inclusion and exclusion criteria. Following this, bidirectional snowballing was conducted utilizing references from the source articles. After that, we created a data extraction form to answer the research questions. A data extraction form was then crafted to address the research questions. Lastly, a data synthesis approach was developed to analyze and summarize the research findings.
Figure 2 depicts the article selection procedure for determining the primary studies used in this SLR. The following subsections provide detailed information about the selection process.
Figure 2: An overview of the steps in determining primary studies
3.1 Research Questions
Our SLR includes all relevant research outlining the various SATD detection techniques. To define the search string, we systematically broke down the research questions using the PICOC framework (Population, Intervention, Comparison, Outcome, and Context), as recommended by Petticrew and Roberts 28. Boolean operators were applied to structure the search effectively: OR was used to include synonyms and alternative terms, while AND linked the key terms, as shown in Table 1. This approach ensured a comprehensive and focused search strategy.
Table 1: PICOC Criteria
Parameter Description Keywords Used
Population Field of application “self-admitted technical debt” OR SATD
Intervention Approach or method for dealing with a certain problem “nlp” AND “technical debt”
Comparison Not aplicable Not aplicable
Outcome Related to factors of importance to practitioners “detection” OR “identification” OR “prediction”
Context Context in which intervention take places “software development” OR “software engineering”
According to the PICOC in Table 1, the main objective of this SLR is to investigate, identify, and classify all relevant evidence from the primary studies that focus on using NLP to detect self-admitted technical debt. Therefore, the aim of this SLR is to answer the following questions:
RQ1: What approaches have been employed to detect self-admitted technical debt (SATD) in software development?
Rationale: This question aims to understand the progress made in detecting SATD. By surveying existing studies, we can detect trends and potential future research directions.
RQ2: What methodologies have been proposed to detect SATD in software development?
Rationale: This question aims to analyze the various approaches and tools developed over the past decade for SATD detection.
RQ3: How effective are these approaches regarding their metrics evaluation results?
Rationale: This question aims to evaluate the performance and practicality of the identified approaches by comparing their precision, recall, and F1-score.
RQ4: What tools have been proposed for SATD detection?
Rationale: This question aims to evaluate the current landscape of tools designed for detecting SATD from the last decade.
RQ5: What are the primary challenges associated with the detection of SATD?
Rationale: This question focuses on uncovering the common obstacles faced when detecting SATD, which could help direct future research and tool development.
3.2 Search Protocol
Before beginning the SLR planning stage, it is necessary to conduct a pilot search on a database such as Scopus to observe if there are any SLRs with the same research theme. There is no need to conduct a new SLR if there are existing SLRs on the same topic 29. In this systematic review of the literature, no results covering the topic of the capabilities of the NLP approach in identifying technical debt were discovered; therefore, the SLR can be conducted as new research.
The SLR’s search strategy consists of three stages: pilot search, primary search, and snowballing. In the pilot search, multiple search strings are used on online databases to determine the most effective one. The primary search uses the identified search string to find relevant literature in online source databases. The search concludes with a snowball search 30. The pilot and primary search utilized six online databases: Web of Science, Scopus, IEEE Xplore, ScienceDirect, ACM Digital Library, and Springer. These databases, as proved by previous software engineering SLRs 31, 32, 33, 34, provide an excellent overview of the software engineering literature and are trustworthy enough to cover our scope of the literature. Moreover, we employed the “Advanced Search” feature in these databases to obtain full-text and metadata results.
Table 2: General search strings utilized during pilot search
String ID Search String
S1 (“self-admitted technical debt" OR SATD) AND (NLP OR “natural language processing”)
S2 (“self-admitted technical debt" OR SATD) AND (NLP OR “natural language processing”) AND (detection OR identification)
S3 (“self-admitted technical debt" OR SATD) OR (“technical debt" AND NLP) AND (detect* OR identif* OR predict*) AND (“software engineering" OR “software development")
The results of a pilot search using search string S1 were communicated to the second author by the first author. The second author selected 5% articles at random to see if they were relevant. The second author proposed altering the search string to S2 because the search string S1 result was still too general. Similar to the first process, iteration is repeated until it is decided to utilize the final search string, namely search string S3.
As illustrated in Table 2, through a pilot search with multiple search strings, the most effective one (i.e., S3) was selected. The search string was designed with synonyms, alternative keywords, and boolean operators ‘OR’ and ‘AND’ to retrieve relevant studies. Drawing from the pilot search conducted earlier, for this search string S3, we incorporated three supplementary terms commonly associated with tasks related to SATD identification: detect, identify, and predict. Additionally, wildcard asterisk (*) was used to broaden the search for variations of words 35.
The final search string (i.e., S3) was applied to the mentioned 6 online databases to find relevant primary studies for the research questions as described in Subsection 3.1. We utilized the “Advanced Search” feature in these databases to acquire full-text and metadata results.
Table 3: Search results from selected databases
No Database Search Result
1 Web of Science 51
2 Scopus 1,004
3 IEEE Xplore 168
4 ACM Digital Library 41
5 Springer 125
6 ScienceDirect 51
Total 1,440
As depicted in Figure 2, the initial step involved searching multiple online databases using predefined search queries to collect a comprehensive set of studies relevant to the research topic. As shown in Table 3), a total of 1,440 studies were retrieved at this stage. To manage such a large dataset, both authors worked collaboratively to remove duplicates, which resulted in 1,119 unique records. Following the duplicate removal, a filtering process based on metadata by using inclusion and exclusion criteria presented in Table 4 was conducted. At this stage, the titles, abstracts, and keywords of the remaining studies were examined to determine their relevance to SATD detection.
A full-text analysis of these 69 studies was then performed to evaluate the quality and relevance of the research in greater detail. The first author initially reviewed the metadata and conducted an independent assessment of the full papers, focusing on their contributions to SATD detection approaches. After completing the initial evaluation, the first author discussed the findings with the second author, who served as a supervisor throughout the process. This discussion provided a critical review of the selection criteria and the individual assessments.
To ensure that no bias influenced the selection process, the second author independently reviewed a subset of the full-text studies. These results were then compared with the first author’s evaluations, and any discrepancies or differences in judgment were carefully discussed and resolved. This collaborative approach minimized the risk of subjective bias and ensured that each study was rigorously and objectively evaluated based on its quality and relevance.
To further ensure consistency in the evaluation process, the test-retest strategy outlined by Kitchenham and Charters 10 was employed, rather than relying on inter-rater agreement. This strategy involved discussions between the first and second authors to maintain a consistent evaluation framework. Additionally, the second author cross-checked the data extraction results, which were shared in a spreadsheet [^1] The spreadsheet is available in the shared dataset area at https://github.com/edisutoyo/satd-detection-slr. Any disagreements that arose during the process were addressed through collaborative discussions until consensus was reached.
After this comprehensive review and resolution of all discrepancies, both authors agreed to select 63 relevant studies from 6 online databases to be included in the systematic review. This thorough, unbiased, and transparent process ensured the inclusion of only the most relevant and high-quality studies.
3.3 Article Selection and Snowballing Process
Inclusion and exclusion criteria must be precisely defined for study selection. The inclusion criteria should include an automated approach to detect technical debt. Consequently, the primary studies were chosen based on comparing titles, keywords, abstracts, and full texts (i.e., introduction, methodology, results, and conclusion). We employed 6 inclusion criteria and 5 exclusion criteria, as shown in Table 4):
Table 4: Inclusion and exclusion criteria
No Inclusion criteria
1 The article must be peer-reviewed and published at a conference, workshop, or journal
2 The article must be accessible online (i.e., articles can be fully accessed and downloaded)
3 For research that has both publication types (journal and proceeding version), only the journal version will be selected
4 The study must discuss approaches for identifying and measuring the performance of SATD detection
5 The study must be published between the years 2014 and 2024
6 The study reported must be related to SATD in software engineering or software development
No Exclusion Criteria
1 Articles written in languages other than English
2 Works in the form of dissertations, theses, and books
3 Articles that are still in the form of proposals
4 The article does not answer at least one research question
5 All other secondary and tertiary studies
To further enhance the comprehensiveness of the review, a bidirectional snowballing technique 30 was employed. Both forward and backward snowballing were performed, meaning that the references cited in the selected studies, as well as the studies that cited them, were examined to identify any additional relevant research that might have been overlooked in the initial search. This process resulted in the inclusion of 5 additional studies. The second author actively contributed to the snowballing process, assessing the relevance and quality of newly identified studies.
In the final stage, the 63 studies selected from the full-text filtering process were combined with the 5 studies identified through bidirectional snowballing, resulting in a total of 68 primary studies. This collaborative effort between the authors ensured that the selection process was both thorough and objective, ultimately leading to a comprehensive review of the existing literature on SATD detection approaches.
3.4 Quality Assessment Criteria
After finalizing the list of articles for the SLR, we performed a quality assessment to ensure that the selected papers provided the necessary information to answer our research questions. The quality of the selected articles was assessed using the questions presented in Table 5, following the procedure recommended by Kitchenham 26. All 68 relevant articles we discovered passed the quality evaluation, and thus, our SLR relies on these 68 primary studies.
Table 5: Quality assessment criteria
No Questions
1 Are the research objectives or research questions in the article clear?
2 Does the study include reviewing the relevant literature or background?
3 Does the paper describe a research methodology?
4 Does the article employ an algorithm or model for SATD detection?
5 Does the study provide research findings and answer the research questions?
6 Does the article report metric evaluation performance or at least report results of the approach?
7 Does the article clearly state the dataset used?
All of the listed articles were published between January 2014 and June 2024. Our searches for articles were completed by the end of June 2024, supplemented by articles obtained through bidirectional snowballing. As a result, it is highly unlikely that we missed any published articles during our period due to publication delays.
3.5 Data Extraction
We selected the final relevant articles after applying the inclusion and exclusion criteria and conducting the quality evaluation. We determined the required information and attributes for data summarization to address our research questions. This facilitated the data extraction process, where we read the full text of each article and extracted the relevant information based on specified attributes. To answer the research questions mentioned in Subsection 3.1, we extracted data items listed in Table 6 from each selected study. The extracted data was then entered into a spreadsheet for further analysis.
Table 6: Data extracted from each study
No Data Name Description Relevant RQ
1 Title Title of the study Overview
2 Author Name of the authors Overview
3 Year The publication year of the study Overview
4 Venue The name of the publication venue of the study Overview
5 Publication Type Journal, conference, workshop, or book chapter Overview
6 Algorithms used Algorithm or model used in the study RQ1, RQ2, RQ4
7 TD Types The types of TD, e.g., design or requirement TD R2, RQ3
8 Number of Class Number of class(es)/TD types used R2, RQ3
9 Dataset used Data source and metadata of dataset used R1, RQ4
10 Metrics Evaluation used Types of metric evaluations used for performance measurement RQ4
11 Performance Evaluations Performance measurement results using metric evaluations RQ4
12 Challenges Issues or challenges related to SATD detection RQ5
4 Results
This section addresses the posed research questions. To that purpose, the outcomes of the 68 primary studies are presented and discussed.
4.1 Results - RQ1: Approaches used for detecting SATD in software development
As part of the results for RQ1, we categorized the relevant studies into three main groups based on their approach: papers using NLP, those employing machine learning (ML), and studies utilizing deep learning (DL) or Transformers (see Figure 3 and Table 7). Any disagreements that arose during the grouping process were resolved through discussion.
Figure 3 shows the number of papers published by approach: NLP, ML, DL, and Transformers. Over the past decade, the identification of SATD has evolved significantly, driven by advancements in NLP and ML techniques. Early methods relied primarily on traditional NLP techniques, focusing on keyword detection and textual patterns in code comments. As depicted in Figure 3, NLP dominated initially, with limited exploration of advanced computational methods. From 2017 to 2019, ML approaches were increasingly introduced to improve SATD detection accuracy. Between 2020 and 2022, the use of DL and Transformer architectures surged, marking a shift toward more sophisticated, data-driven techniques. These models captured complex patterns in software artifacts, improving SATD detection. As of 2023 and 2024, DL and Transformer models continue to lead SATD detection efforts.
Figure 3: Articles count throughout the years based on their approaches
Table 7: Approach identified. Studies that fall into more than one approach are highlighted with square brackets
Approach Paper Reference Technique examples
NLP No. 8, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49 POS tagging, Regular Expressions, Textual patterns, etc.
Machine Learning No. 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64 65, 66, 67, 68, 69, 70, 71, 72,[73, 74, 75, 76] NB, ME, SVM, kNN, etc.
Deep Learning No. 77, 78, 79, 80, 81, 82, 82, 83, 84, 85, 86, 87, 88, 89 90, 91, 92, 93, 94,[73, 74, 75, 76] CNN, LSTM, BiLSTM, etc.
Transformers No. 95, 96, 97, 98, 99, 100,[73, 74, 75] BERT, ALBERT, GPT, etc.
NLP approaches have played a fundamental role in detecting SATD. These methods generally concentrate on text preprocessing tasks that encompass stemming, lemmatization, and stopwords removal. The objective is to process and evaluate textual data from code comments or documentation, where technical debt is frequently acknowledged by the authors themselves. As presented in Table 7, other methods include manual analysis and Parts-of-Speech (POS) tagging 36, 40, 41, 49, which aid in classifying and identifying relevant information. Regular expressions 38, 46, keyword-labeled 61, and textual patterns 45 are commonly employed to semi-automate the identification of certain technical debt indicators in large codebases.
As the complexity of software systems increased, so has the demand for more sophisticated approaches in technical debt identification, resulting in the use of ML classifiers. These algorithms include Naïve Bayes (NB) 50, 55, 59, Maximum Entropy (ME) 51, 67, Support Vector Machines (SVM) 52, and k-Nearest Neighbor (kNN). Logistic Regression (LR) 101 and Random Forest (RF) 56, 43 are also employed, with various degrees of accuracy and interpretability in detecting patterns linked with technical debt. Furthermore, XGBoost 66, 68, a relatively new addition to the ML classifier, improves performance by refining gradient boosting approaches.
The rise of DL has introduced even more advanced techniques for detecting SATD. Convolutional Neural Network (CNN) 77, 83, 85, 90, Long Short-Term Memory (LSTM) 79, 81, and Bidirectional Long Short-Term Memory (BiLSTM) 78, 82 are among the most prominent techniques used. They are particularly useful in processing structured data and identifying patterns that are not easily captured by traditional ML algorithms. Deep learning approaches have been extensively studied and are shown to be more critical in automatically detecting technical debt.
Transformers are the most recent development in the detection of SATD. By offering excellent context understanding and language representation, these models, namely Bidirectional Encoder Representations from Transformers (BERT) 95, 96, 100, 97, A Lite BERT (ALBERT) 98, Robustly Optimized BERT (RoBERTa) 98, and DistilBERT 98, have transformed the field of NLP. Transformers excel at handling large-scale data and understanding the nuances in code comments and other textual software artifacts that signify technical debt. The literature reflects a growing trend towards using these models to achieve higher accuracy and more robust identification processes, especially in dealing with large and complex datasets typical of modern software systems.
Answer to RQ1 Four main approaches have been identified to detecting SATD: traditional NLP, ML, DL, and Transformers. Over time, the focus has shifted from basic NLP techniques to more advanced methods like ML and DL, with transformers now leading the way, due to their effectiveness in capturing complex textual patterns.
4.2 Results - RQ2: Proposed methodologies for detecting SATD in software development
To address this, we categorized the methodologies into three main types: manual, semi-automated, and automated. These categories represent the evolution in improving SATD detection’s accuracy, efficiency, and scalability. Manual methods depend on human expertise to review text artifacts, while semi-automated techniques combine human input with tools to streamline the process. Fully automated methods use advanced algorithms to detect SATD with minimal human intervention. Each approach has its own strengths and challenges, which we will explore. Disagreements in categorization were resolved through discussion.
Figure 4: Articles count throughout the years based on the methodologies
Over the past decade, methods for detecting Self-Admitted Technical Debt (SATD) have changed significantly, as shown in Figure 4. There has been a clear shift from manual to more automated techniques, driven by a growing need for efficiency and accuracy in managing technical debt in software development.
From 2014 to 2016, SATD detection mainly relied on manual analyses. This process involved developers or analysts closely examining codebases and documentation to spot technical debt. While these manual methods provided valuable insights, they were time-consuming and susceptible to human error, highlighting the demand for more efficient solutions.
Between 2017 and 2019, the field began transitioning to semi-automated techniques. These methods combined manual efforts with automated tools to enhance the SATD detection process. The figure illustrates a gradual increase in the adoption of semi-automated approaches during this time, often utilizing regular expressions or Part-of-Speech (POS) tagging to reduce developer workload and improve consistency.
The period from 2020 to 2022 marked a significant move towards fully automated SATD detection methods. As depicted in the figure, there was a notable rise in the use of automated techniques, indicating the maturation of the field. These approaches employ advanced technologies such as natural language processing (NLP), machine learning (ML), and deep learning (DL) to automatically detect technical debt with minimal human input. As we move into 2023 and 2024, the trend toward automation in SATD detection continues to grow. The figure shows that automated methods have become the primary approach, outpacing manual and semi-automated techniques. This suggests that the future of SATD detection will likely focus on refining automated tools, possibly integrating more advanced AI and ML models.
As presented in Table 8, the methodologies are classified into three broad categories: Manual, Semi-Automated, and Automated approaches. These categories are determined based on the extent of human intervention and the utilization of NLP, ML, DL, Transformers models.
Table 8: Methodologies identified
Methodology Paper Reference NLP ML DL TRANSFORMERS
Manual No. 8, 37, 59, 39, 102 No. 8, 37, 59, 39, 102 - - -
Semi-Automated No. 36, 38, 40, 61, 41, 44, 45, 65, 46, 47, 49, 48 No. 36, 38, 40, 41, 44, 45, 46, 47, 49, 48 No. 65, 61 - -
Automated No. 50, 51, 52, 53, 54, 55, 56, 57, 77, 58, 79, 60, 62, 42, 43, 63, 64, 80, 81, 82, 83, 74, 84, 101, 85, 86, 87, 66, 88, 67, 89, 73, 68, 69, 76, 90, 91, 92, 70, 93, 94, 75, 95, 96, 99, 100, 71, 72, 97, 98 - No. 50, 51, 52, 53, 54, 55, 56, 57, 58, 60, 61, 42, 43, 64, 84, 101, 66, 67, 73, 68, 69, 70, 71, 72 No. 77, 78, 79, 62, 63, 80, 81, 82, 83, 74, 85, 86, 87, 88, 89, 76, 90, 91, 92, 94, 96 No. 93, 75, 95, 99, 100, 97, 98
Manual approaches significantly depend on human intervention, usually requiring the manual annotation of code comments or other textual software artifacts to detect SATD. Manual methods solely rely on NLP techniques. There is no application of ML, DL, or Transformers in the manual category, as these techniques are generally used for more automated processes. These methods are frequently time-consuming but offer high precision in identifying subtle or context-specific technical debt items. For example, Potdar and Shihab 8 conducted research using source code comments from several open-source projects. They manually classified comments as SATD or non-SATD based on specific comment patterns. Similarly, Maldonado and Shihab 37 employed manual labeling to construct a dataset for SATD that focused on several specific types of debt, including design debt, requirement debt, documentation debt, test debt, and defect debt. This strategy was critical in establishing baseline datasets that could subsequently be utilized to train automated classifiers 51.
Semi-automated approaches combine manual review with automated tools to improve efficiency while retaining accuracy. Semi-automated methods predominantly rely on NLP, including part-of-speech (POS) tagging 36, and textual patterns 38; and a few of them use ML approaches 65, 61. However, there is no indication of the usage of Deep Learning or Transformers in these semi-automated techniques.
Automated approaches are characterized by the use of ML, DL, or Transformers and advanced NLP techniques to fully automate the SATD detection process. In order to detect SATD from issue tracking systems, Dai and Kruchten 50 created an automated method that makes use of an ML algorithm, namely of a Naïve Bayes (NB) classifier in addition to other NLP approaches like TF-IDF and TextRank. This method demonstrated the potential for high scalability and integration into software development processes, with promising results in terms of precision, recall, and F1-score. Recently, Li et al. 76 introduced a novel approach to automatically detect SATD from four different sources: source code comments, commit messages, pull requests, and issue tracking systems. The most recent trend in automated SATD detection includes the use of Transformers-based models (e.g., BERT 93, 75, RoBERTa 98, 95, and GPT 99).
Table 9: Textual artifacts Used. Studies that use more than one textual software artifact are highlighted with square brackets
Artifact Paper Reference NLP ML DL TRANSFORMERS
Code comment No. 8, 36, 37, 38, 51, 52, 53, 54, 55, 56, 57, 77, 58, 78, 40, 40, 60, 61, 62, 41, 42, 63, 64, 80, 81, 44, 82, 83, 45, 84, 101, 85, 86, 65, 66, 88, 67, 89, 73, 46, 68, 91, 92, 70, 93, 94, 75, 96, 99, 71, 49, 48, 72, 97,[69, 76, 90, 100] 8, 36, 37, 38, 40, 41, 44, 45, 46, 49, 48 51, 52, 53, 54, 55, 56, 57, 58, 60, 61, 42, 64, 84, 101, 65, 66, 67, 73, 68,[69], 70, 71, 72 77, 78, 79, 62, 63, 80, 81, 82, 83, 86, 88, 89,[76], [90], 91, 92, 94, 96 93, 75, 99,[100], 97
Issue No. 50, 59, 39, 43, 102, 74, 87, 95, [69, 76, 90, 100] 59, 39, 102, 45 50, 84,[69] 74, 87,[76, 90] 95,[100]
Pull request No. 98,[69, 76, 100] - [69] [76] 98,[100]
Commit message No. 47,[69, 76, 100] 47 [69] [76] [100]
Moving forward from the methodological categorization, it is essential to understand the source or type of textual artifacts analyzed in the context of SATD. Table 9 categorizes the various studies on SATD detection based on the types of textual artifacts utilized. Code comments (such as from Java projects, blockchain projects, Dockerfile, and XML files), issues (from issue tracking systems and GitHub issue reports), pull requests, and commit messages contain valuable information for SATD detection, providing varying levels of granularity and relevance depending on the artifact analyzed. The application of NLP, ML, DL, and Transformers-based models across these artifacts shows how methodologies have evolved to handle different levels of complexity and scale.
Code comments are the most commonly analyzed artifact for SATD detection, with NLP, ML, DL, and Transformers models being widely employed (e.g., studies 8, 36, 93). While issues, pull requests, and commit messages are also explored, they are less frequently studied. Early research on SATD primarily relied on NLP to extract patterns from developer language in code comments and issues. However, more recent studies have incorporated ML, DL, and Transformers models (e.g., 76, 93, 98) for a deeper and more automated analysis, especially when dealing with larger and more complex datasets. The increasing adoption of Transformers models reflects the field’s shift towards more sophisticated and scalable SATD detection across various textual artifacts.
Meanwhile, there has been a noticeable shift towards using a more diverse set of textual artifacts for detecting SATD. Furthermore, studies that utilize more than one textual software artifact (e.g., code comments combined with issue trackers or commit messages) are highlighted with square brackets. The increased use of pull requests and commit messages (e.g., 69, 76, 100) as sources of SATD detection in recent studies signals a growing recognition of these artifacts’ value in providing temporal and contextual data on technical debt. As the field moves forward, there is an increasing emphasis on automation and leveraging diverse textual artifacts, which aims to improve the scalability and reliability of SATD detection across varied software projects. The studies that employed multi-artifacts represent an effort to enhance the accuracy and robustness of SATD detection by cross-referencing different sources of developer commentary and interaction.
Answer to RQ2 Over the past decade, there has been a visible shift from manual towards fully automated approaches to improve efficiency and scalability in SATD detection. Additionally, the use of multiple textual artifacts has grown, enhancing the accuracy and robustness of SATD detection by integrating diverse sources from the software development process.
4.3 Results - RQ3: Effectiveness of the proposed approaches regarding their metrics evaluation results
To evaluate the effectiveness of the proposed SATD detection approaches, we examined the performance metrics reported in the studies included in our review. While a subset of the studies did not provide detailed performance results, the majority (49 out of 68) reported on key evaluation metrics such as precision, recall, and F1-score. The following section presents a detailed analysis of these metrics, utilizing boxplots to visualize the performance variability and trends across different SATD detection approaches.
Out of 68 relevant studies, 17 did not report their performance results, and 2 only provided AUC scores. Therefore, 49 out of the 68 articles included in our review reported performance results using at least one of the common evaluation metrics: precision, recall, or F1-score. We collected all performance results presented in these 49 articles and organized them into a spreadsheet. Rather than focusing solely on the best-performing model from each study, we included the results from all model classifiers. This approach provides a more comprehensive and unbiased evaluation of SATD detection models 103. By considering all models, researchers can capture a broader range of performance variability, which is essential for understanding the factors that influence model effectiveness. Limiting the analysis to only the best-performing models could lead to skewed conclusion, as it overlooks the context in which the models perform and fails to account for suboptimal models that offer important insights into potential limitations or challenges. Moreover, analyzing all models enables a more robust cross-study comparison, allowing for the identification of consistent trends and patterns in model performance across different contexts, datasets, and techniques. This holistic approach strengthens the reliability of the conclusion and ensures their applicability across a wider range of scenarios, rather than being limited to cases of optimal performance.
We use boxplots to demonstrate the effectiveness of various approaches, as they are a useful graphical tool for showing variations between populations without relying on assumptions about data distribution 103. We compiled the effectiveness of various techniques as reported in the original papers, utilizing precision, recall, and F1-score derived from calculations based on the confusion matrix. Through boxplots, we showcase studies on precision, recall, and F1-score in relation to different factors of the models.
  • Precision = $\frac{TP}{TP+FP}$
  • Recall = $\frac{TP}{TP+FN}$
  • F1-score = $\frac{2*Precision*Recall}{Precision+Recall}=\frac{2*TP}{2*TP+FP+FN}$
We acknowledge the potential for misinterpretation when aggregating the results of approaches across various studies without considering the diversity of datasets utilized. The rationale behind this methodological choice is to cultivate an inclusive understanding of each model’s performance and versatility across a wide range of contexts and scenarios. This inclusive approach enables the identification of universally effective strategies as well as those that excel under specific conditions, thereby providing a nuanced perspective on the potential applicability, hereby offering a holistic view of their effectiveness in detecting SATD. This methodology aligns with precedents in the field, notably reflected in the systematic literature review, such as the study of Hall et al. 103 and Malhotra 104. This precedent underscores the validity and relevance of cross-dataset comparisons in extracting meaningful insights that transcend the particularities of individual studies. By synthesizing findings across diverse datasets, this review aspires to contribute to the advancement of SATD detection approaches, offering valuable guidance for future research and practical applications in software engineering.
To avoid any misleading interpretations, we divide the process of SATD identification into two distinct tasks: SATD identification and SATD categorization. SATD identification focuses on determining whether a particular entity or instance is present in the text of a software development artifact, classifying it as either SATD or non-SATD. SATD categorization, on the other hand, takes this a step further by not only identifying the presence of SATD but also classifying it into specific types of technical debt. These categories may include labels such as code debt, design debt, requirement debt, and others, based on the linguistic elements found in the text.
In the following subsection, we present performance results for each SATD detection task, illustrated through boxplots.
4.3.1 Performance results of SATD identification
The box plot shown in Figure 5 provides a comparative overview of performance across different SATD identification methods, specifically for binary classification, which aims to accurately distinguish between SATD and non-SATD instances. The three key metrics displayed in the box plot—precision, recall, and F1-score—are standard measures for evaluating the effectiveness of classification models. Precision indicates the model’s ability to correctly identify instances as SATD or non-SATD, while recall measures how effectively the model captures all relevant SATD instances. The F1-score, which represents the harmonic mean of precision and recall, offers a balanced evaluation, especially when one metric is prioritized over the other. Together, these metrics provide valuable insights into the performance of SATD detection methodologies in identifying technical debt within software projects.
Figure 5: Boxplot of SATD identification grouped by their approaches
The NLP-based approach demonstrates a relatively high precision of 0.79, indicating that it successfully identifies SATD in a significant number of cases when it classifies an instance as SATD. However, its recall is lower at 0.62, suggesting that it overlooks a considerable portion of SATD instances. The F1-score of 0.66 reflects the imbalance between precision and recall, highlighting that while NLP methods were foundational in early SATD detection, they now face challenges in fully capturing all relevant SATD instances.
In contrast, Machine Learning (ML) models offer a more balanced performance, achieving a precision of 0.73 and recall of 0.74, which results in an F1-score of 0.73. This improvement over NLP methods illustrates ML’s effectiveness in reducing false positives while maintaining good coverage of SATD instances. Historically, ML approaches have represented a significant advancement in SATD detection, effectively bridging the gap between precision and recall.
Deep Learning (DL) models further enhance performance, with precision and recall scores of 0.78 and 0.76, respectively, yielding a balanced F1-score of 0.76. This suggests that DL techniques are better at identifying SATD instances while minimizing false positives. The improvements in both precision and recall demonstrate the advancements deep learning architectures have brought to the complexities of SATD identification, enabling more accurate classification across diverse software contexts.
The Transformers-based approach achieves the best overall performance, with a precision of 0.83—the highest among all methods—and a recall of 0.76, leading to an F1-score of 0.76. These results underscore the superior capability of Transformers models to accurately identify SATD while retrieving a majority of relevant instances. As the latest advancement in SATD detection techniques, Transformers represent the current state-of-the-art, providing both the precision needed to avoid false positives and the recall required to capture a wide range of SATD instances.
Referring to Figure 5, we can further elaborate and illustrate the metric results in more detail, organizing them based on their respective models, as depicted in Figure 6. Each model is evaluated with a specific number of experiments (denoted by “n="), highlighting the range of experiments used to measure their performance.
NLP-based models, particularly Textual Patterns and Annotation Tags, were assessed for their effectiveness. Textual Patterns (n=71) achieved relatively high precision, indicating that this model accurately classifies SATD examples with low false positives. However, its lower recall (0.62) and F1 score (0.66) suggest that it often fails to identify many SATD instances, limiting its overall effectiveness in comprehensive identification. In contrast, Annotation Tags (n=21) performed better in recall and F1 score, indicating a more balanced detection capability, although variability was present, as shown by the wider interquartile range.
ML models exhibited a broader range of performance. Among them, Random Forest (RF) (n=34) and Naive Bayes (NB) (n=37) provided the most consistent results, achieving high precision and F1 scores. This indicates their effectiveness in reducing false positives while capturing a majority of SATD instances. However, models like Support Vector Machine (SVM) (n=13) and k-NN (n=10) showed lower precision and recall, indicating challenges in accurately identifying SATD instances. Notably, Auto-sklearn and LightGBM performed well, though their small sample sizes (n=3 and n=1, respectively) limit the generalizability of their results.
DL models have made significant strides in SATD identification. BiLSTM (n=168), CNN (n=151), and LSTM (n=100) achieved high scores across all metrics, demonstrating consistent precision, recall, and F1 scores. These models effectively handle complex SATD detection tasks while minimizing false positives. However, models like Graph Neural Network (GNN) (n=35), Recurrent Neural Network (RNN) (n=20), and Graph Convolutional Network (GCN) (n=20) exhibited lower performance, particularly in recall, indicating they may miss important SATD instances. Models with smaller samples, such as Ladder Network (LN) (n=10) and Gated Graph Neural Network (GGNN) (n=10), also showed variability in performance, likely due to their limited test sizes.
Transformer-based models represent the most recent advancements in SATD identification. Among these, BERT (n=69), RoBERTa (n=39), and Transformer (n=40) achieved the highest precision, recall, and F1 scores. These models demonstrated remarkable consistency and reliability in identifying SATD instances, with minimal variability, making them leading techniques in SATD detection. Other Transformer models, such as ALBERT (n=18), DeBERTa (n=10), and GPT (n=3), also performed well, though slightly less robustly than the BERT-based models. DistilBERT (n=2) showed promise, but its limited sample size makes it difficult to draw definitive conclusions about its overall performance.
Figure 6: Models performance of SATD identification grouped by each approach
4.3.2 Performance result SATD categorization
Figure 7 offers insight into the precision, recall, and F1-score of these approaches, focused on various SATD categorization methodologies for multiclass classification, and provides insight into how effectively these compiled approaches differentiate between types like design debt, requirement debt, documentation debt, test debt, defect debt, and others. The metrics provide insights into how effectively these approaches perform in a multiclass classification scenario, where distinguishing between different categories of SATD is more complex than a binary classification.
Figure 7: Boxplot of precision, recall, and F1-score of SATD categorization
NLP approaches demonstrate a low Recall of 0.11 and an F1-score of 0.16, indicating significant difficulty in capturing true SATD instances and missing many relevant examples. This limitation affects their effectiveness in real-world applications, particularly in complex multiclass classification tasks. In contrast, ML approaches show notable improvement, achieving a Precision of 0.75 and a Recall of 0.57, resulting in a balanced F1-score of 0.52. These results suggest that ML models provide better accuracy in SATD classification, although they still miss a considerable number of true positive cases, as indicated by the Recall.
DL approaches perform well across all metrics, with a strong Precision of 0.77 and a high Recall of 0.76. However, the F1-score of 0.62 suggests that while they successfully capture most true positives, there is still potential for improvement in balancing false positives and false negatives. Interestingly, Transformer-based approaches, despite being more advanced, exhibit slightly lower overall performance in this instance, with a Precision of 0.69, Recall of 0.59, and an F1-score of 0.54. While transformers typically excel in complex text classification tasks, these results may reflect the influence of model configuration or data characteristics on their performance in this specific SATD identification task. Nonetheless, transformers still outperform NLP approaches and are comparable to ML models in this context.
Referring to Figure 7, we can provide a comprehensive evaluation of specific types of SATD categorization, organized by their respective models, as shown in Figure 8. Each model is evaluated with a specific number of experiments (denoted by “n="), highlighting the range of experiments used to measure their performance.
Figure 8: Models performance of SATD categorization grouped by each approach
In the NLP group, two approaches, Textual patterns (n=46) and Annotation tags (n=2), were evaluated. The Textual Patterns approach showed highly variable performance, with precision hovering around 0.6, while the recall and F1 scores dropped significantly, indicating that while this method can accurately categorize SATDs in some cases, it struggles to consistently retrieve all SATD instances. The large interquartile range and the presence of outliers highlight the inconsistency of this method in different contexts. In contrast, Annotation tags, although evaluated on only two samples, showed relatively stable and high precision, recall, and F1 scores, indicating strong performance for the multiple cases tested, albeit with limited generalizability. These findings confirm the limitations of NLP techniques in handling the complexities of SATD categorization. While effective in some cases, particularly for high precision, NLP methods often miss critical SATD instances, leading to suboptimal recall and overall lower F1-scores. This aligns with the review’s conclusion that NLP approaches, although foundational in early SATD detection, are no longer sufficient for comprehensive and consistent categorization of SATD.
ML models showed a wide range of results, reflecting the diversity in algorithm complexity and adaptability to SATD categorization. Random Forest (RF) (n=114) and Auto-sklearn (n=108) showed the highest performance, consistently providing high precision and recall, indicating that these models are able to balance false positives and missed SATD examples effectively. However, Naive Bayes (n=17) showed weaker performance, particularly in terms of recall, indicating that it may miss important SATD cases. Other models, such as SVM (n=25), XGBoost (n=50), and Logistic Regression (n=24), showed lower recall and more significant performance variation across the dataset, reflecting their challenges in categorizing SATD consistently.
DL models showed significant improvement over the NLP and ML approaches, specifically in their ability to maintain high precision, recall, and F1 scores across larger data sets. CNN (n=163), CNN-BiLSTM (n=20), and LSTM (n=70) stood out for their consistent and robust performance, with higher recall rates indicating their effectiveness in detecting a wider range of SATD events without significantly increasing false positives. GGNN and Feedforward neural network (FFNN), despite their smaller sample sizes, also showed strong performance, although the small dataset makes it difficult to draw broader conclusion about their generalizability.
Transformers-based models, specifically BERT (n=47), ALBERT (n=24), and RoBERTa (n=13), were the highest performing approaches in SATD categorization. BERT emerged as the best performer, with high precision, recall, and F1 scores across all evaluations, demonstrating its ability to accurately and consistently detect SATDs with minimal false positives and missed events. ALBERT and RoBERTa also performed strongly, although evaluated on smaller sample sizes, their high precision and F1 scores reinforced their value in SATD detection.
From Figures 6 and 8, we can determine the best-performing algorithm for each group based on precision, recall, and F1-score. It is important to note that Figure 6 represents SATD identification aimed at distinguishing SATD from non-SATD (binary classification), while Figure 8 illustrates the results of SATD categorization of specific SATD types (multiclass classification). To select the best-performing algorithm from the boxplot, we primarily focused on the F1-score (dark blue bars) as it offers a balance between precision and recall. We identified the algorithm with the highest median F1-score, indicated by the middle line of the boxplot, and considered its variance (the spread of the box and whiskers), with lower variance suggesting more consistent performance. Additionally, we examined the precision and recall (light and mid-blue bars) to ensure a balance between the two, prioritizing algorithms with high values in both. In addition, we minimized the influence of algorithms with numerous outliers, as they indicate inconsistent performance, even if their median scores were high. Finally, we factored in the n value (number of experiments), giving preference to algorithms tested more frequently, as a higher n suggests more reliable and stable results across different trials.
Table 10: Summary of highest performance results based on SATD detection tasks
Approach Metric Evaluations
Precision Recall F1-score
SATD identification
NLP Annotation tags Annotation tags Annotation tags
ML SMO LR SMO
DL GNN BiLSTM BiLSTM
Transformers GPT Transformer BERT
SATD categorization
NLP Annotation tags Annotation tags Annotation tags
ML SMO NB SMO
DL GGNN GGNN GGNN
Transformers BERT BERT BERT
The best-performing algorithms for each approach based on the different SATD types is presented in Table 10. In both the SATD identification and categorization tasks, BERT consistently emerges as the best-performing model. For SATD identification, BERT’s high median F1-score, tested across a large number of experiments (n=69), showcases its strong reliability and generalizability. Its low variance and minimal outliers further demonstrate its consistent and stable performance across different datasets. Leveraging its advanced Transformer-based architecture, BERT excels at capturing the context and nuances in SATD instances, making it highly effective in balancing precision and recall.
Similarly, in the SATD categorization task, BERT outperformed other models, achieving the high median F1-score with a significant number of experiments (n=47). Its minimal variance and absence of outliers reinforce its reliability and consistency in this task as well. Altogether, these factors position BERT as the most robust and effective model for both SATD identification and categorization, making it the top choice in each task.
Answer to RQ3 BERT (a Transformers-based model) emerges as the top performer in both SATD and non-SATD identification and the categorization of specific SATD types, providing balanced precision, recall, and F1 scores. SMO (an ML model) and GGNN (a DL model) also show strong performance within their respective domains, with SMO excelling in precision and F1 score, while GGNN leads in identifying specific SATD types. Overall, Transformer models, particularly BERT, represent the most advanced and effective approach to SATD detection.
4.4 Results - RQ4: Proposed tools for SATD detection
To address this research question, we conducted a thorough review of existing tools developed over the past decade. While the detection and management of SATD have been extensively studied, there remains a noticeable gap in translating research into practical tools for developers. This section examines tools specifically designed for SATD detection, offering insights into advancements in automation and machine learning, as well as the diverse approaches used to address various forms of technical debt.
Table 11 presents a collection of tools proposed for SATD identification across different software projects. This compilation reflects the evolution of tool development over time, highlighting an increasing emphasis on automation, machine learning integration, and varied strategies for managing different types of technical debt. Below, we describe these tools and their significance in the broader context of SATD identification, emphasizing key trends that have emerged over the last decade.
Table 11: SATD detection tools proposed
Name Reference Year Task Type Description Status
eXcomment No. 36, 48 2015 Identification A tool designed to parse Java source code and fetch code comments to identify SATD or non-SATD Broken link
SATD Detector No. 55 2018 Identification A Java library and an Eclipse plug-in. The tool automatically identifies whether comments are SATD or non-SATD and integrates with an IDE to help manage these comments Obsolete
DebtHunter No. 64 2021 Identification, Categorization A machine learning-based approach for detecting self-admitted technical debt Available
FixMe No. 72 2021 Categorization A GitHub bot that is developed to detect, monitor, and notify developers about On-hold SATD in their repositories Broken link
SATDBailiff No. 67 2022 Identification A tool designed to mine, identify, and track SATD Available
A browser extension No. 74 2022 Categorization A browser extension utilizes a developed ML model to automatically categorize the SATD types (i.e., documentation, code, defect, test, design, build, architecture, versioning, usability, and requirement debt) of a rOpenSci R package Not available
DebtViz No. 90 2023 Categorization A tool for SATD detection, classification, visualization, and monitoring in a single platform to categorize code/design, documentation, test, and requirement debt types Available
The tools proposed for SATD identification reflect a diverse range of approaches to addressing technical debt in software projects. eXcomment 36, 48, and SATD Detector 55 focus on identifying both SATD and non-SATD comments, with eXcomment integrating into an IDE for real-time detection. A browser extension 74 leverages machine learning to detect various SATD types, including documentation and usability debt, highlighting the potential of ML for broader SATD detection. SATDBailiff 67 provides comprehensive tracking and monitoring of SATD over time, while FixMe 72 is a GitHub bot specifically designed to alert developers about unresolved, “on-hold SATD" in repositories. Lastly, DebtViz 90 offers a more complete solution, combining SATD detection, classification, visualization, and monitoring in one platform, emphasizing the need for visual tools that make technical debt more understandable to stakeholders. Together, these tools demonstrate an increasing focus on automation, monitoring, and comprehensive SATD management, though the overall number of tools remains limited relative to the research interest in the field.
This trend of limited tool development is further highlighted when considering that, despite the SLR encompassing 68 relevant studies, only 7 articles proposed specific tools for SATD identification. This disparity points to a concerning gap between the significant research interest in SATD and the actual development of practical, automated solutions that can assist developers in managing technical debt. While a range of tools has been introduced, the overall lack of available tools indicates that much of the research has yet to translate into actionable, industry-ready solutions. Addressing this gap will be critical for advancing the practical management of technical debt in software projects, moving from theoretical frameworks toward robust, scalable automation.
Answer to RQ4 We identify several tools proposed for SATD detection, with a focus on automation, machine learning integration, and addressing different types of technical debt. However, despite significant research interest, only a few tools have been developed, revealing a gap between theoretical advancements and the availability of practical, automated solutions for managing technical debt in real-world software projects.
4.5 Results - RQ5: Primary challenges associated with SATD detection
To address the challenges associated with SATD detection, we conducted a thorough analysis of the relevant literature, focusing on the difficulties commonly highlighted by researchers. These challenges reflect the complexities inherent in identifying SATD across various software projects and methodologies.
We identified nine key challenges related to SATD identification extracted from the relevant studies, as summarized in Table 12, with relevant studies emphasizing each challenge. These challenges were determined by examining sections commonly located in the ‘Discussion’ or ‘Threat to Validity’ sections of the papers reviewed. We extracted all discernible challenges, whether implicitly or explicitly mentioned, associated with SATD detection. Each challenge is explained in this section and further elaborated on below.
Table 12: Papers highlighting key challenges. Studies that fall into more than one challenge are highlighted with square brackets
No Challenge Paper Reference
1 Error-proneness dues to manual analysis No. 8, 38,[37, 36, 50, 53, 59, 39, 49]
2 Imbalanced data No. [54, 77, 78, 40, 79, 60, 61, 62, 64, 81, 82, 83, 45, 85, 86, 88, 68, 93, 94, 96, 71], 56, 63, 80, 66, 89, 46, 47, 92, 75, 95, 97
3 Generalizability No. [50, 53, 59, 39, 37, 55, 77, 40, 60, 61, 64, 78, 81, 45, 85, 86, 88, 68, 93, 94, 96, 49, 99, 71], 58, 74, 101, 87, 65, 69
4 Explainability No. [77, 78, 83, 85, 86, 88, 99, 44, 100]
5 False negatives No. [54, 79, 62]
6 Risk of reproducibility No. [ 44, 100, 67]
7 Computational costs No. [44, 82, 83, 67, 100]
8 Overlap of features among debt types No. [37, 66]
9 Tagging and comment relevance No. [36, 49]
Manual Analysis
The first challenge is the dependence on manual analysis, which introduces the potential for human error and inconsistency 38. Since the process relies on human judgment, it becomes prone to inconsistencies in labeling, as different individuals may interpret and categorize SATD differently 36. This subjectivity in analysis can lead to varying outcomes in tasks like pattern recognition and feature engineering, making the process less reliable.
The variability in human interpretation introduces a lack of uniformity in the labeled datasets, which can severely impact the effectiveness of machine learning models that rely on this data for training. When models are trained on datasets that contain inaccuracies or inconsistencies, their predictive accuracy and generalization capabilities suffer. This challenge becomes even more pronounced when scaling up to larger datasets, where consistent and precise labeling is critical to maintaining the quality of the training data. Without robust and reliable analysis, the models may struggle to handle the increased complexity of larger data, leading to reduced performance and misidentification of SATD instances.
Imbalanced data
In most software projects, SATD comments represent only a small portion of the total code comments, resulting in a significant class imbalance. This imbalance can adversely affect the performance of SATD identification models, as the models may tend to focus on non-SATD comments, which dominate the dataset. As a result, the models may struggle to accurately identify the relatively rare SATD instances, often prioritizing the majority class and overlooking the minority class 66.
This imbalance frequently leads to a high number of false negatives 105, where the model fails to detect SATD instances, undermining its effectiveness. Since machine learning models generally find it difficult to learn patterns from underrepresented classes 106, the scarcity of SATD examples makes it challenging for these models to develop a robust understanding of SATD characteristics. Consequently, the model’s ability to identify and classify SATD instances diminishes, particularly when faced with a large volume of non-SATD items. Addressing this imbalance is critical to improving the precision and reliability of SATD identification models, as it ensures that the models are not biased toward detecting the majority class while neglecting the crucial SATD instances.
Generalizability
One of the primary challenges associated with the identification of SATD is the issue of generalizability. Generalizability refers to the ability of SATD identification models to maintain their accuracy and effectiveness when applied across different software projects and domains. This challenge arises due to the diverse nature of software projects, which often exhibit significant variations in coding practices, domain-specific terminologies, and architectural designs 44.
Models trained on data from one specific project tend to be tailored to the characteristics of that project, limiting their performance when applied to other projects with different practices. For example, the way SATD is expressed in comments may differ across projects due to differences in the development team’s language, domain jargon, or code style conventions 58, 39. These project-specific factors can result in models that are highly effective within their training environment but struggle to generalize to other contexts, especially when faced with unfamiliar coding practices or new domains.
Explainability
Another primary challenge in the SATD identification process, particularly in machine learning and deep learning-based models, is the lack of explainability. This issue arises because many advanced models, such as neural networks and deep learning architectures, function as “black boxes," meaning their internal decision-making processes are not easily interpretable by users 107. These models generate predictions without providing transparent explanations about the underlying reasoning for classifying an instance as SATD, non-SATD, or even specific types of SATD.
The absence of explainability presents a significant obstacle to the adoption and trustworthiness of SATD detection tools. Moreover, explainability is crucial for validating the accuracy of SATD identification models. If developers cannot trace the logic behind a model’s predictions, it becomes difficult to diagnose errors or refine the model to reduce false positives or false negatives.
False negatives
False negatives are another concern in SATD detection. They occur when SATD items are incorrectly classified as non-SATD, leading to undetected technical debt in the codebase. This happens when models fail to identify instances where developers have either implicitly or explicitly admitted technical debt 38. As a result, undetected technical debt can linger in the system, leading to long-term code degradation, higher maintenance costs, and an increased risk of introducing more bugs and system failures 12, 108.
Risk of reproducibility
Another primary challenge in the context of machine learning and supervised models-based models is the issue of reproducibility 44, 67. These models often involve complex architectures with numerous parameters that must be fine-tuned for optimal performance. However, many studies fail to provide sufficient details on the hyperparameters, training configurations, or the specific processes used to achieve their results. Additionally, the lack of publicly available code exacerbates the reproducibility problem, as other researchers are unable to replicate the exact experimental setup. This makes it difficult to reproduce results across different studies and datasets, raising concerns about the reliability and validity of these approaches.
Computational costs
Another challenge associated with SATD detection is the high computational costs of many advanced models, particularly deep learning approaches. These models require significant resources during training, including high-performance GPUs, large memory capacities, and extended computation time. This high computational cost creates several obstacles for researchers and practitioners who aim to implement or improve SATD identification methods.
Additionally, the extended computation time required for training deep learning models presents a challenge in terms of practicality. SATD identification models often need to be retrained or fine-tuned to adapt to new projects or datasets, and the time-consuming nature of this process can hinder agile software development practices. In scenarios where rapid feedback is essential, such as continuous integration and deployment environments, the lengthy training times of deep learning models can become a bottleneck, delaying the detection of technical debt and the subsequent mitigation efforts.
Overlap of features among debt types
Furthermore, the overlap of features among different types of technical debt, such as code debt, design debt, and requirement debt, presents another challenge 66. This overlap makes distinguishing between different debt types difficult, leading to potential misclassification. The ambiguity in distinguishing between SATD and other debt types can reduce the accuracy of identification models. If features that are common to multiple debt types are not carefully handled, models may struggle to correctly identify SATD.
Tagging and comment relevance
Finally, tagging and comment relevance is a critical challenge in identifying SATD. Developers often use tags such as “TODO, “FIXME," or “XXX" in code comments or other textual artifacts to mark areas that require attention or indicate temporary fixes. While these task-related tags are commonly employed as indicators of SATD, they are not always relevant to technical debt. In many cases, tags like “TODO" or “FIXME" can represent a simple reminder or a non-technical debt issue, thereby creating noise in the identification process. Moreover, not all SATD instances are explicitly marked by such tags, further complicating the identification process.
Answer to RQ5 We identified nine major challenges in detecting SATD, including issues like manual analysis errors, imbalanced data, lack of generalizability, and explainability in advanced models. Other challenges include false negatives, reproducibility risks, computational costs, overlapping features among debt types, and the relevance of tagging and comments, all of which hinder effective SATD identification.
5 Discussion
The findings from this systematic review reveal several key trends and challenges in the detection of SATD over the past decade, demonstrating both the evolution and fragmentation within the field. SATD detection approaches have progressed significantly, shifting from simple keyword-based NLP methods to more complex ML, DL, and Transformers-based models. While each advancement has brought increased performance and automation, the development of practical and scalable SATD detection tools remains limited.
5.1 Evolution of SATD detection techniques
Early SATD detection techniques primarily relied on basic NLP methods, such as keyword matching and textual patterns, to identify technical debt in code comments. These approaches, although foundational, were often limited by their inability to account for the nuanced language used by developers and their reliance on rule-based systems that could not easily adapt to different contexts. From 2017 onward, ML techniques such as Naive Bayes (NB), Support Vector Machines (SVM), and Random Forests became more prevalent, offering improved accuracy by learning from labeled datasets and reducing false positives. Despite this progress, these methods still faced challenges related to generalizability across different datasets and projects, as well as difficulty in explaining the rationale behind the classification of certain comments as SATD.
The rise of deep learning methods from 2020 to 2022 marked a significant leap forward in SATD detection. Models like Long Short-Term Memory (LSTM) networks and Convolutional Neural Networks (CNN) enabled the identification of more complex patterns in developer comments and project artifacts. These approaches improved recall and reduced false negatives, offering a more balanced approach to identifying SATD across diverse datasets. However, the high computational cost of training and deploying these models remains a major barrier to widespread adoption.
Most recently, Transformers-based models, such as BERT and RoBERTa, have emerged as the leading approaches for SATD detection. These models have demonstrated superior performance in capturing the context and semantics of developer comments, enabling more accurate and generalizable SATD identification. The ability of Transformers to handle large-scale datasets and complex text has pushed the boundaries of what is possible in SATD detection. However, issues related to the explainability of these models and their scalability for industrial use still persist.
5.2 Challenges in SATD detection
Despite the advancements in SATD detection techniques, several challenges hinder the practical adoption of these methods in real-world software development environments. A key challenge in SATD detection is the imbalance in datasets—SATD instances are often much less frequent than non-SATD instances, which negatively impacts model performance. This imbalance is not limited to the distinction between SATD and non-SATD but also occurs among specific SATD types, where certain categories, such as requirements debt or design debt, are far less represented, further complicating detection. Although several techniques have been employed to address data imbalance in SATD detection, future research should explore more advanced methods, such as leveraging large language models (LLMs) to better handle imbalanced data and improve model performance.
Generalizability remains another key challenge. Many SATD detection models are trained on specific datasets, often from open-source projects, which limits their ability to perform well on new or proprietary codebases. This lack of generalizability reduces the practical utility of these models in diverse software development environments, where coding practices, project documentation, and domain-specific language can vary significantly. The explainability of advanced models, particularly deep learning and Transformer-based approaches, is also a concern. Without a clear understanding of why certain comments are classified as SATD, developers may be reluctant to trust and adopt these models in their workflows. Future work should focus on developing explainable AI techniques for SATD detection to ensure that these models’ decision-making processes are transparent and understandable to developers. Research on enhancing the interpretability of DL and Transformers models, perhaps through visualizations or simplified intermediate representations, would help practitioners better trust and utilize these models in real-world software development settings.
Another significant barrier to adoption is the high computational cost associated with training and maintaining deep learning and Transformers-based models. While these models offer improved accuracy, their resource-intensive nature makes them difficult to implement for smaller development teams or in environments with limited computational resources. This challenge highlights the need for more efficient algorithms or alternative approaches that can deliver high accuracy without the associated computational burden.
5.3 Lack of tools for SATD detection
Despite the significant academic interest in SATD detection, the translation of research into practical tools remains limited. This review identified only seven tools developed specifically for SATD detection over the past decade, reflecting a notable gap between theoretical research and practical application. While tools like SATD Detector and eXcomment provide valuable functionality for identifying SATD in Java codebases, the overall number of available tools is insufficient to meet the needs of diverse software development teams. Additionally, the tools that do exist often lack the ability to integrate seamlessly into modern development environments, limiting their utility for real-time SATD management.
The limited development of practical tools also highlights the need for further collaboration between academia and industry. Researchers must focus on creating tools that not only demonstrate high performance in academic settings but also address the specific needs and constraints of industrial software development, such as scalability, usability, and integration with existing development environments. Bridging the gap between theoretical advancements and practical, scalable tools is essential for achieving widespread adoption in industry. Future research should prioritize developing user-friendly tools that can be integrated into continuous integration (CI) environments, enabling real-time SATD detection and providing actionable insights for developers. Moreover, these tools should facilitate collaboration within teams working on large-scale software projects, ensuring that they are both practical and effective in real-world development scenarios.
5.4 Future Directions
To address the challenges identified in this review, future research should prioritize enhancing the generalizability, explainability, and efficiency of SATD detection models. Techniques such as transfer learning, which allow models to be trained on one dataset and then applied to others with minimal retraining, could significantly improve generalizability. At the same time, incorporating explainable AI (XAI) techniques is crucial to ensure that the decision-making processes of these models are transparent and understandable for developers. Developing more interpretable models, or employing post-hoc explanation techniques, could foster trust in these tools and encourage their adoption in real-world projects. Further research on enhancing the interpretability of deep learning and Transformer models—potentially through visualizations or simplified intermediate representations—would empower practitioners to better understand, trust, and effectively utilize these models in software development settings.
There is also a need for the development of more practical SATD detection tools that can be easily integrated into existing development environments. These tools should focus on providing real-time feedback to developers, enabling them to address technical debt as part of their normal workflow. By bridging the gap between research and practice, the software engineering community can ensure that SATD detection techniques contribute meaningfully to the long-term maintenance and evolution of software systems.
6 Threats to Validity
As with any systematic literature review (SLR), the results of this study are subject to various validity threats. These are categorized as internal, external, and construct validity threats, along with mitigation strategies to minimize their impact.
6.1 Internal validity
Internal validity refers to the reliability of the methodology employed in the study. One key threat to internal validity in this SLR is related to the formulation of the search strings. In the pilot search, several iterations of search strings were tested, starting with a more general search (e.g., “self-admitted technical debt" OR SATD and NLP), which led to a broader, less focused set of results. Over time, more refined strings were developed, including terms like “detection" and “identification," as well as the use of boolean operators and the wildcard (*) for broader term variations.
The final search string (S3), which combined terms such as “self-admitted technical debt" OR SATD with NLP OR “natural language processing" AND (detection OR identification OR prediction), was chosen after multiple pilot tests to ensure it captured the most relevant studies. This iterative process helped mitigate the risk of missing key studies during the initial literature search.
Furthermore, data extraction inaccuracies could threaten the reliability of our findings. To reduce this risk, the data extraction process was double-checked, and disagreements were resolved through discussions between the authors. A pilot data extraction was also conducted to further minimize potential bias.
6.2 External validity
External validity refers to the extent to which the findings of this SLR can be generalized to other contexts beyond the studies reviewed. A primary threat to external validity in this research is the limited generalizability of the results due to the scope of the reviewed datasets. All of the studies included in this review rely on open-source projects, which may not fully capture the complexities or constraints faced in proprietary or large-scale industrial environments.
Open-source datasets often have characteristics, such as well-documented codebases or community-driven development practices, that differ significantly from proprietary software projects, which may involve stricter deadlines, unique domain requirements, or higher security and compliance constraints. As a result, some of the trends and findings related to SATD detection may not extend to these more complex and regulated environments. For example, the performance of machine learning or deep learning models, which may perform well on open-source datasets, could degrade when applied to more diverse or proprietary datasets with varying development practices.
6.3 Construct validity
Construct validity pertains to whether the study accurately measures the concepts it aims to investigate. A significant threat in this SLR is the comparison of results from different studies, which can be misleading if the studies use different methodologies or report metrics inconsistently. Directly comparing the precision, recall, and F1-scores of models trained on heterogeneous datasets without considering their context could lead to inaccurate conclusion.
To address this, we employed a grouping strategy to compare similar studies and ensured that the datasets were organized before any comparisons were made. By contextualizing the results within groups, we aimed to make the comparisons more meaningful. Furthermore, this methodology is based on established practices within the field, notably exemplified by Hall et al. 103 and Malhotra 104, which similarly undertakes the task of comparing results across different studies to deduce broad-based insights into algorithmic performance in varying contexts.
7 Conclusion
This SLR provides an in-depth analysis of SATD detection approaches over the past decade, shedding light on the evolution of techniques from traditional keyword-based methods to sophisticated machine learning and deep learning models. Among the key findings, it is evident that while NLP-based techniques were foundational in early SATD detection, they struggle with recall and overall performance in comparison to ML and DL approaches. Transformer-based models, particularly BERT, emerged as the most effective in capturing complex patterns in textual artifacts, offering superior performance across multiple metrics.
However, several challenges persist, including issues related to dataset diversity, the scalability of models, and the explainability of AI techniques. The reliance on open-source datasets in many studies limits the generalizability of the findings, and more research is needed to adapt these models to varied industrial environments. Moreover, the imbalance of SATD datasets continues to affect the robustness of model performance, and integrating multiple textual artifacts such as commit messages and pull requests presents further complexity. Furthermore, the lack of widely adopted tools for SATD detection in industry highlights a gap between research advancements and practical applications.
In conclusion, while the field of SATD detection has advanced considerably, there remains a need for further research and collaboration between academia and industry to ensure these advancements translate into practical tools that enhance software maintainability and reduce the long-term impact of technical debt.
[…]
*Data availability Supplementary material related to this article, including all the boxplots as images, can be found online at https://github.com/edisutoyo/satd-detection-slr
[…]
*Declaration of competing interest The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
[…]
*Acknowledgement This work was financially supported by the Indonesian Education Scholarship (BPI) from the Center for Financing of Higher Education (BPPT) and Indonesia Endowment Fund for Education (LPDP).